Stock Y has a beta of 0.8 and an expected return of 8.71 percent. Stock Z has a beta of 1.9 and an expected return of 13.36 percent. What would the risk-free rate (in percent) have to be for the two stocks to be correctly priced relative to each other? Answer to two decimals.
Answer #1

Expected return (CAPM) = Rf+β×Rp

Rf is risk free return

Rp is risk premium

Rf+0.8×Rp = 8.71% ...... (1)

Rf+1.9×Rp = 13.36% …… (2)

Do (1)-(2)

1.1×Rp = 4.65%

Market risk premium, Rp = 4.23%

Take, Rf+0.8×Rp = 8.71%

Rf+0.8×4.23% = 8.71%

Risk free rate of return, Rf = 5.33%

*Please rate thumbs up